OpenStudy (anonymous):

How do I find the center and radius of a circle for the equation x^2 +(y-1)^2=4

OpenStudy (asnaseer):

do you know the general form for the equation of a circle?

OpenStudy (asnaseer):

if not, this might help: http://www.mathsisfun.com/algebra/circle-equations.html

OpenStudy (anonymous):

thank you again :-)

OpenStudy (asnaseer):

yw :)

OpenStudy (anonymous):

It helped but I am still confused again ): I see the formulas but I am not sure which one to use or how to apply the equation to the formulas..

OpenStudy (asnaseer):

ok, if a circle has its center at (a, b), then what would its equation be?

OpenStudy (anonymous):

maybe it would help if you rewrote your equation this way: \(\large (x-0)^2 +(y-1)^2=4 \)

OpenStudy (asnaseer):

and if it had a radius r

OpenStudy (anonymous):

\[(x-a)^2 + (y-b)^2=r^2\]

OpenStudy (anonymous):

?

OpenStudy (asnaseer):

good - so now try and rewrite your equation in this form

OpenStudy (anonymous):

\[x^2 +y^2 -2y-3=0 \]

OpenStudy (anonymous):

I ended up with this...

OpenStudy (asnaseer):

no - do not expand the terms

OpenStudy (asnaseer):

your original equation is:\[x^2 +(y-1)^2=4\]so the term in y is already in the "right" form (i.e. \((y-b)^2\))

OpenStudy (asnaseer):

now think of how you can rewrite \(x^2\) as \((x-a)^2\) - what would a=?

OpenStudy (anonymous):

4?

OpenStudy (asnaseer):

so you are saying that \(x^2\) is the same as \((x-4)^2\)?

OpenStudy (anonymous):

no, i was moving the 4 over and somehow making that work. i am not sure how i can rewrite it.

OpenStudy (asnaseer):

maybe this will help:\[x^2=(x)^2=(x-?)^2\]what should the ? be replaced with to make this correct?

OpenStudy (anonymous):

1?

OpenStudy (anonymous):

ohhh 0?

OpenStudy (asnaseer):

which means you are saying that:\[x^2=(x-1)^2\]try again...

OpenStudy (asnaseer):

bingo! well done!

OpenStudy (anonymous):

lol it takes me awhile to catch on, thank you for being patient with me.

OpenStudy (asnaseer):

so now we have:\[(x-0)^2+(y-1)^2=4\]but we need it in the form:\[(x-a)^2+(y-b)^2=r^2\]so think about how you can express 4 as something squared?

OpenStudy (anonymous):

two squared, correct?

OpenStudy (asnaseer):

perfect! so we ended up with:\[(x-0^2+(y-1)^2=2^2\]so you should now be able to tell me where the center of this circle is and what its radius is

OpenStudy (asnaseer):

sorry - missed a bracket on the x terms there

OpenStudy (anonymous):

the center is (0,1) and the radius is 2?

OpenStudy (asnaseer):

perfect!

OpenStudy (anonymous):

thank you :-)

OpenStudy (asnaseer):

yw :)

OpenStudy (anonymous):

I also had to find the intercepts for that equation. I graphed the circle and found 4 intercepts, but two of them were wrong. My answer was (2,0), (0,3),(-2,0), and (0,-1). The correct answer is \[(-\sqrt{3},0),(\sqrt{3},0), (0,-1), (0,3)\]

OpenStudy (anonymous):

what did I do wrong?

OpenStudy (asnaseer):

you do not need to graph this to find the intercept. use the same method as I showed you before to find the intercepts - i.e. set x=0 to find the y-intercepts and set y=0 to find the x-intercepts.

OpenStudy (anonymous):

oops lol thanks

OpenStudy (asnaseer):

BTW: I guess since you are fairly new to this site you should know that the usual workflow is: 1. You post a question. 2. One or more people come to your question to help you out. 3. You select the one that best answered your question by clicking on the blue Best Response button next to any one of their replies. 4. You close the question by clicking on the Close button then the whole process starts again.

OpenStudy (anonymous):

I'm sorry

OpenStudy (asnaseer):

np :)